# Constraining string cosmology with the gravitational-wave background using the NANOGrav 15-year data set

### v1

@@ 1 Introduction
Despite these remarkable achievements, the precise origin of the observed PTA signals remains uncertain [15, 16], with hypotheses encompassing both astrophysical and cosmological sources. The diverse range of potential origins includes supermassive black hole binaries (SMBHBs) [17, 18, 19, 20, 21], as well as more exotic phenomena such as scalar-induced GWs [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40], cosmic phase transitions [41, 42, 43, 44, 45, 46, 47, 48, 49], domain walls [50, 51, 52], cosmic strings [53, 54, 55, 56, 57, 58], and modified gravities [59, 60, 61, 62, 63, 64, 65]. Another intriguing possibility is that the detected signal may be the SGWB originating from string cosmology.

### v2

@@ 1 Introduction
Despite these remarkable achievements, the precise origin of the observed PTA signals remains uncertain [15, 16], with hypotheses encompassing both astrophysical and cosmological sources. The diverse range of potential origins includes supermassive black hole binaries (SMBHBs) [17, 18, 19, 20, 21], as well as more exotic phenomena such as scalar-induced GWs [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], cosmic phase transitions [44, 45, 46, 47, 48, 49, 50, 51, 52], domain walls [53, 54, 55, 56], cosmic strings [57, 58, 59, 60, 61, 62], and modified gravities [63, 64, 65, 66, 67, 68, 69]. Another intriguing possibility is that the detected signal may be the SGWB originating from string cosmology.